# A Survey on Tabular Data Generation: Utility, Alignment, Fidelity, Privacy, and Beyond

### v1

@@ 6.2 Methods
GAN-based models. TableGAN: Another model that marked an important milestone for deep generative modelling of tabular data is TableGAN [?], which, unlike medGAN, can model continuous features. While it is based on a standard GAN model, consisting of a generator and a discriminator, TableGAN has an additional component: a classifier that predicts the value of the target feature for synthetic samples using semantic integrity constraints learned from the real data. As exemplified in the original paper, a sample with values of $60$ and $1$ for features Cholesterol and Diabetes, respectively, does not align with the background knowledge, as the cholesterol level is too low for the target feature to indicate a diabetes diagnosis. Focusing on ensuring privacy, the authors argue for the use of synthesiser models, which do not learn bijective relationships between real and fake data, making them less vulnerable to re-identification attacks, unlike prior perturbation or anonymisation methods. Indeed, TableGAN shows that it can synthesise tabular data with a low risk of re-identification, membership and attribute disclosure attacks to succeed. PATE-GAN: Another model that is based on a GAN architecture, but alters it for preserving privacy, is PATE-GAN [?]. Rather than using the standard GAN discriminator, PATE-GAN relies on a modified Private Aggregation of Teacher Ensemble (PATE) [?] mechanism to ensure that the discriminator is differentially private. More specifically, a number of teacher-discriminators are created and trained to improve their loss w.r.t. the generator, which adjusts its loss according to the single student-discriminator in the mechanism. In turn, the student-discriminator adjusts its loss according to the teacher-discriminators and allows for backpropagation to the generator, closing the loop. Critically to the privacy requirement, specialising each of the teacher-discriminators on small partitions of the training set ensures that the effect of individual samples is small and, thus, higher differential privacy. CuTS: ? [?] proposed CuTS as a customisable framework, which can adapt to user-defined specifications, such as background knowledge constraints, but also to differentially private training. In the private setting, CuTS builds on the DP iterative framework from [?], but uses its own GAN-based generator, which is equipped with a mechanism that matches the generator’s output distribution to the real data distribution by comparing marginals. Additionally, CuTS can incorporate background knowledge in the form of propositional logical constraints that each data point must satisfy via a new loss term. Further, rejection sampling is used to ensure that the model’s outputs satisfy the constraints. Although the framework works only with discrete features, it takes a step towards synthesisers for tabular data that account for multiple requirements, including alignment and privacy.
7 Other models
@@
In (§3-6), we compared deep generative models originally proposed for tabular data synthesis. Here, we discuss seminal models first developed in other fields and later adapted for tabular data. WGAN: Originally designed for computer vision tasks, WGAN [?] was introduced in a paper that analyses ways to measure the distance between the real and model distributions, arguing that weaker loss functions induce weaker model topologies, thus weak models for the real distribution. The authors compare the Wasserstein distance with traditional probability distance measures, including Kullback-Leibler (KL) divergence, Jensen-Shannon divergence, and total variation distance, showing that the Wasserstein distance has properties that would be better suited for learning distributions in low dimensional manifolds. They thus propose GAN models that use the Wasserstein distance as loss function. WGAN-GP: A popular extension of WGAN is WGAN-GP [?], which added a new term in the loss that penalised the gradient norm of the discriminator w.r.t. the discriminator’s input, providing an alternative to the weight clipping that was leading to the generator retaining the real data points and adding Gaussian noise to fill the gaps. VEEGAN: VEEGAN [?] jointly trains a generator and a novel reconstructor network in order to avoid mode collapse. The reconstructor learns to map the real data distribution to Gaussian random noise, approximating the inverse of the generator and encouraging the generator to map the noise distribution back to the real data distribution.
@@ 8 Discussion and Future Directions
In this survey, we identified four key requirements for the effective deployment of synthetic data, along with the relevant evaluation procedures and deep generative models, at the same time categorising the latter by their model architecture.
@@
Related work. Previous surveys generally compare methods by the architecture-specific properties of the models. For example, both ? (?) and ? (?) group methods mainly by their architectures, while also describing the necessary data transformation and regularisation techniques. [?], provide an overview of the models proposed in the field and the metrics used to evaluate them, while not taking this requirement-centric perspective which helps both newcomers and seasoned practitioners and researchers in the field to find the perfect model and evaluation protocol for their needs. Further, they exclude diffusion-based and LLM-based architectures.
Future Directions. Tabular data generation is a rapidly evolving field still in its early stages. As illustrated in Figure 2, many early approaches have prioritised individual objectives—typically utility or privacy—leading to highly specialised models that lack broad applicability. This fragmentation highlights a key challenge: the need for models that can balance multiple, often competing, requirements.
It is worth to be noted that alignment has historically received limited attention but is now emerging as a crucial factor for generating realistic data. With the increasing deployment of generative models in real-world applications and the growing prominence of neurosymbolic AI and safe AI [??], we anticipate that alignment will soon become a fundamental requirement in the field.
Moreover, Figure 2 reveals a striking gap: no existing model successfully integrates all the core requirements we consider. As the field matures, we expect new requirements will emerge, and the hybridisation of techniques will drive the development of more versatile models. Bridging these gaps will be essential for advancing the field and enabling broader, more reliable applications of tabular data generation.
Acknowledgments

### v2

@@ 4.1.1 GAN-based models
TableGAN: Another model that marked an important milestone for deep generative modelling of tabular data is TableGAN [Park et al., 2018], which, unlike medGAN, can model continuous features. While it is based on a standard GAN model, consisting of a generator and a discriminator, TableGAN has an additional component: a classifier that predicts the value of the target feature for synthetic samples using semantic integrity constraints learned from the real data. As exemplified in the original paper, a sample with values of $60$ and $1$ for features Cholesterol and Diabetes, respectively, does not align with the background knowledge, as the cholesterol level is too low for the target feature to indicate a diabetes diagnosis. Focusing on ensuring privacy, the authors argue for the use of synthesiser models, which do not learn bijective relationships between real and fake data, making them less vulnerable to re-identification attacks, unlike prior perturbation or anonymisation methods. Indeed, TableGAN shows that it can synthesise tabular data with a low risk of re-identification attacks, membership inference attacks and attribute disclosure attacks to succeed. PATE-GAN: Another model that is based on a GAN architecture, but alters it for preserving privacy, is PATE-GAN [Yoon et al., 2019]. Rather than using the standard GAN discriminator, PATE-GAN relies on a modified Private Aggregation of Teacher Ensemble (PATE) [Papernot et al., 2017] mechanism to ensure that the discriminator is differentially private. More specifically, a number of teacher-discriminators are created and trained to improve their loss w.r.t. the generator, which adjusts its loss according to the single student-discriminator in the mechanism. In turn, the student-discriminator adjusts its loss according to the teacher-discriminators and allows for backpropagation to the generator, closing the loop. Critically to the privacy requirement, specialising each of the teacher-discriminators on small partitions of the training set ensures that the effect of individual samples is small and, thus, higher differential privacy.
CuTS: Vero et al. [2024] proposed CuTS as a customisable framework, which can adapt to user-defined specifications, such as background knowledge constraints, but also to differentially private training. In the private setting, CuTS builds on the DP iterative framework from [McKenna et al., 2022], but uses its own GAN-based generator, which is equipped with a mechanism that matches the generator’s output distribution to the real data distribution by comparing marginals. Additionally, CuTS can incorporate background knowledge in the form of propositional logical constraints that each data point must satisfy via a new loss term. Further, rejection sampling is used to ensure that the model’s outputs satisfy the constraints. Although the framework works only with discrete features, it takes a step towards synthesisers for tabular data that account for multiple requirements, including alignment and privacy.
4.1.2 GAN and BN-based models
GANBLR: Zhang et al. (2021) construct the generator and discriminator as Bayesian networks (BNs), which are able to encode known feature interactions. The background knowledge inclusion leads to better utility than standard GANs. C3-TGAN: Similarly to GANBLR, C3-TGAN [Han et al., 2023] uses Bayesian networks to capture background knowledge. However, while GANBLR models only discrete features, C3-TGAN can handle both discrete and continuous features by following the CTGAN [Xu et al., 2019] approach. To incorporate explicit attribute correlations and property constraints from background knowledge, it represents constraints as control vectors (similar to the conditional vectors in CTGAN) and uses them to guide training, ensuring better alignment.
4.1.3 GAN and AE-based models
ITS-GAN: proposed for tabular data augmentation ITS-GAN [Chen et al., 2019] is an early method advocating for the use of background knowledge. The knowledge it can express is of two types: (i) rules stating that the value of one feature uniquely determines the value of another (e.g., the feature Position determines the feature Salary), and (ii) rules stating that a specific value for a set of features determines the specific value for another (e.g., if Position = “CEO" then Salary = “2M"). To encode the first types of rules, they train on the real data an autoencoder (AE) for each rule to output the value of the dependent features given the values of the independent ones. Then, at training time, the GAN generator is trained to minimise the difference between its output and one of the AEs, while also being penalised if it violates any rule of the second type. Additionally, the discriminator receives as input the difference between the AE’s outputs and the synthetic samples, thus being able to use this information to assess the sample.
medGAN: Proposed by Choi et al. [2017] for synthesising healthcare datasets, medGAN is one of the pioneering models for synthesising tabular data using deep generative models. medGAN generates distributed representations of health records using the generator of a GAN-based model, then decodes these representations using an autoencoder pretrained on the real data, and passes the decoded representations to the discriminator component of the GAN, which is trained to distinguish the synthetic from the real data. In terms of privacy, medGAN showed low attribute disclosure risks, except when an attacker focused on a small number of data points. IT-GAN: Similar to OCT-GAN but designed to enhance privacy robustness, IT-GAN [Lee et al., 2021] employs a generator based on neural ODEs to balance utility and privacy protection. Instead of directly synthesising samples in the input space, the generator produces hidden representations, as seen also in medGAN. Notably, it ensures that the hidden representation space matches the input dimensionality, enabling the application of ODEs.
4.1.4 GAN- and Transformer-based models
TabTransGAN [Zhang et al., 2025] achieves a performance comparable to GReaT [Borisov et al., 2023] by combining the strengths of Transformer [Vaswani et al., 2017] and GAN-based models, which allows it to model both contextual and structural relationships across all features. Specifically, TabTransGAN adopts a GAN architecture in which the generator is based on CTGAN, while the discriminator replaces the standard design with a Transformer architecture. This enables it to better capture both feature-wise distributions and dependencies between the features, leading to more accurate discrimination between synthetic and real data.
4.2 Diffusion-based Methods
4.2.1 Diffusion-based models
TabDDPM: Motivated by the their success in computer vision, Kotelnikov et al. 2023 introduced TabDDPM: a diffusion-based model for synthesising tabular data. Popular for its high utility results, TabDDPM is able to model both discrete and continuous features, but separately, by using multinomial [Hoogeboom et al., 2021] and Gaussian diffusion models [Sohl-Dickstein et al., 2015], respectively. For a simple abstraction on how a diffusion model can be used to generate data, see the central schema in Figure 2. STaSy: Proposed by Kim et al. [2023], STaSy is based on a score-based generative model (SGM), which was shown to be a competitive alternative to diffusion-based models in [Ho et al., 2020]. Contrarily to TabDDPM, StaSy estimates the score function (gradient of log-probability) instead of explicitly modelling the forward and reverse Markov processes. As the loss is difficult to train, STaSy uses a self-paced learning strategy, which starts with a subset of training data—yielding a stable, low loss—and gradually expands to the full dataset. Evaluated on 15 real-world datasets against 7 baselines (mostly GAN-based models), STaSy achieves high utility. Due to the forward and backward passes, both StaSy and TabDDPM suffer from slow sampling times. Moreover, the base SGM model is known to be unstable in high-dimensional settings. CoDi: Motivated by the difficulty of modelling discrete features, Lee et al. [2023] proposed CoDi, a framework comprising co-evolving diffusion models that separately handle continuous and discrete features as in TabDDPM [Kotelnikov et al., 2023], with the difference that the two models condition on each other during training. At each training step, the discrete (resp., continuous) diffusion model takes the perturbed sample from the continuous (resp., discrete) one as input, and both models are conditioned on both samples from the previous step during the denoising process. Slower than most GAN-based models (excluding OCT-GAN), CoDi is faster than STaSy and TabDDPM, while also obtaining higher utility on mixed types datasets.
TDGGD: The first framework encoding background knowledge about feature relations into diffusion-based models is TDGGD [Jia et al., 2024]. The relations can express lower and upper bounds over single features or their sum. Rather than explicitly modelling the relations, TDGGD models whether the features are part of such relations or not, which only gives an indication of hidden relations between the columns to the model. FinDiff: Motivated by the need to preserve the statistical properties of real data of mixed types in real-world financial contexts, Sattarov et al. [2023] proposes FinDiff, which is capable of generating mixed-type tabular financial data using a diffusion model to capture high-dimensional dependencies. Unlike another popular diffusion model, TabDDPM [Kotelnikov et al., 2023], which uses one-hot encoding for categorical features, FinDiff employs an embedding encoding for better handling of mixed data types. This leads to higher fidelity performance compared to the baselines, indicating that FinDiff effectively captures both feature-wise and joint distributions of the data.
4.2.2 Diffusion and AE-based models
AutoDiff: Proposed in [Suh et al., 2023], AutoDiff is a diffusion-based model designed to preserve statistical fidelity by modelling the joint distribution via an autoencoder, rather than modelling the distributions of individual features separately, as seen in GAN- and diffusion-based models like CTGAN and TabDDPM. More precisely, AutoDiff (i) uses the autoencoder to learn the distribution of the features (which can be of mixed types) in a continuous space, and then (ii) passes these continuous representations to the diffusion model, which generates latent representations that are then decoded back into representations in the original feature space. Moreover, like CTAB-GAN, it introduces an approach to handle mixed-type features by adding a new feature to the autoencoder that encodes the frequency of recurring values in a mixed-type feature.
4.2.3 Diffusion and VAE-based models
TabSyn: Introduced by Zhang et al. [2024], TabSyn trains a score-based diffusion model, like STaSy, but in a joint VAE-learned space of numerical and categorical features. Notably, TabSyn significantly reduces the runtime of diffusion-based models while outperforming baselines w.r.t. fidelity and utility.
4.3 Transformer and Large Language Model-based Methods
GReaT [Borisov et al., 2023] represents an answer to the abundance of works that convert tabular data into numerical representations thus losing the contextual connections between the features, which very often carry semantic meaning. Indeed, GReaT performs the following steps (also illustrated in Figure 2): first, the tabular data undergoes a textual encoding process, converting it into text. Next, a feature order permutation step is applied. The resulting sentences are then used to fine-tune the large language models (LLMs) [Vaswani et al., 2017]. At sampling time, either a single feature name or arbitrary feature-value pairs are given as input to the LLM, which then completes them (note that this allows for arbitrary conditioning). Using this approach, GReaT outperforms most of the baselines w.r.t. utility, but its data sampling time is notably high. To address this problem, TabuLa [Zhao et al., 2025] introduces a method to reduce the length of the generated token sequences. Unlike GReaT, which relies on large language models pre-trained on natural language processing tasks, TabuLa starts with a randomly initialised model and iteratively fine-tunes it on tabular data synthesis tasks. Another key distinction is that TabuLa does not use GReaT’s feature permutation mechanism. Instead, it targets scenarios where arbitrary feature conditioning is not needed and assumes that feature values appear in a consistent order across all data points. To this end, the authors tokenise the dataset feature-wise, ensuring that each generated token can be mapped to a specific feature based on its absolute position in the row. Specifically, for each feature, they determine the longest token sequence across the dataset and pad all sequences (corresponding to that feature) to this length during training, Using this approach, TabuLa manages to significantly reduce the training time compared to GReaT. However, its sample generation time is still considerably higher than GAN-based models such as CTGAN or TableGAN.
4.3.1 Tabular Foundational Models
TabPFN Bringing a shift in tabular data synthesis, away from training synthesisers on each new dataset, the tabular prior-data fitted network (TabPFN) Hollmann et al. [2025] leverages in-context learning [Brown et al., 2020] mechanism of transformer-based methods and is pre-trained on a large collection of synthetic datasets rather on than large collections of publicly available datasets. In this way, TabPFN avoids issues such as privacy breaches, copyright infringements, and cross-contaminating the training data with test data. To generate synthetic datasets, it uses structural causal models, which are able to represent causal relationships in the data via a directed acyclic graph. Although primarily a predictive model building on its preliminary version from [Hollmann et al., 2023], it is important to note that TabPFN can also generate tabular data by approximating the joint distribution of features and targets, and is able to outperform the classical methods, such as tree-based baselines, on small- or even medium-sized datasets of up to 10K samples. Given its transformer-based architecture, sampling time during inference is high, but the model intrinsically does not require training, as it can adapt and do inference on unseen real-world datasets.
4.4 Variational Autoencoder-based Methods
4.4.1 VAE-based models
TVAE [Xu et al., 2019], proposed along with CTGAN, was designed to check whether CTGAN’s generator’s lack of access to the real data during training makes it weaker compared to models that do use these data for training their generator, such as models based on Variational Autoencoders (VAEs) [Kingma & Welling, 2014]. Indeed, TVAE’s architecture showed an improvement over CTGAN in terms of fidelity. However, like CTAB-GAN+, the authors also considered the impact of their proposed models w.r.t. privacy and noted that TVAE’s access to the training data might make it unsuitable in downstream tasks where sensitive data are present.
4.4.2 VAE and GNN-based models
GOGGLE: [Liu et al., 2022] is a different approach which integrates knowledge about pairwise feature dependencies by encoding them in a graph where each feature is a node and an edge exists if a relation between the two features is known. New relations can be learnt using a message passing mechanism from graph neural networks (GNNs), which is jointly trained with a VAE-based architecture, where only relationships prioritised by the learned graph influence feature generation. Due to its relational learning graph-based component, it might be difficult to scale GOGGLE to datasets of higher dimensionality.
4.5 Constraints-enforcing Methods
C-DGM: The first method to constrain deep generative models and guarantee the constraints satisfaction was proposed in [Stoian et al., 2024]. The constraints can capture any set of linear inequalities over the feature space. Differently from the methods surveyed above, this approach relies on a layer to be added right before the sample output layer, which restricts the output space of the model to coincide with the space defined by the constraints. The layer can be added on top of any deep generative model (DGM), and the resulting models are called C-DGMs. For a visual representation on how the layer-based methods can be added on a GAN model, see the right-most schema in Figure 2. As the layer is differentiable and acyclic, it can be added both at inference and training time, while also improving the utility of the model. DGM+DRL: The above layer was further extended in  [Stoian & Giunchiglia, 2025], in order to capture constraints as expressive as disjunctions over linear inequalities (which allow for expressing relations like “if the sum of two features is greater than 10 then the difference of other two features should be lower than 5”), thus modelling non-convex and even disconnected output spaces. The new layer is called Disjunctive Refinement Layer (DRL), and a DGM constructed using DRL is called DGM+DRL. Just as C-DGM, DGM+DRL is able to guarantee alignment with the background knowledge, while also improving utility across all considered baselines.
4.6 Other Deep Learning Methods
Normalising flows-based models. Neural Spline Flows: Rather than using the common transformations found in normalising flows, Durkan et al. [2019] proposed using a fully-differentiable module based on monotonic rational-quadratic piecewise functions. This adaptation, combined with the inherent properties of normalising flows, which model the data as the output of an invertible differentiable transformation of a noisy sample drawn from a known distribution, can help synthesise high-fidelity tabular data. The authors note that such a result is conditional on having enough training data compared to the datasets dimensionality.
Energy-based models. TabPFGen [Ma et al., 2023] is a generative model that uses a pretrained network as an energy-based model for data augmentation and class balancing. Although its capabilities are limited to low-dimensional inputs (it was tested on datasets with maximum 10 features), TabPFGen demonstrates good utility performance in the data augmentation task by simply using a pretrained model compared to specialised GAN-, VAE-, and diffusion-based models.
Iterative privacy-preserving NN models. GEM: Addressing the computational bottleneck of differentially private query release methods such as MWEM [Hardt et al., 2012], Liu et al. [2021] propose new generative networks with the exponential mechanism (GEM), which can be incrementally updated without retraining from scratch. The novelty of this approach is its capability to represent the data distribution using a set of generative models parametrised by neural networks. Unlike the prior methods, this approach avoids the issues which arise from maintaining a joint distribution over the data domain as the number of features increases, all while being able to generate privacy-preserving datasets. GEM+: While GEM offers better scalability than the prior methods from the literature, it cannot scale to the real-world high dimensionality demands and also suffers from lower fidelity compared to the more recent methods, such as AIM [McKenna et al., 2022]. Conversely, AIM offers high fidelity but fails to scale to high dimensions due to the memory constraints of graphical models. To tackle this issue, GEM+ [Maddock et al., 2025] proposes a novel model which integrates the adaptive selection strategy of AIM with GEM’s scalable generative networks, while also bringing algorithmic improvements specific to high-dimensional data.
Kernel-based NN Models. DP-MERF: Part of the line of frameworks that develop differentially private algorithms for training deep generative models is the private mean embeddings with random features (DP-MERF) method proposed by Harder et al. [2021]. Specifically, DP-MERF uses random Fourier features to approximate the kernel, which in turn allows for releasing a differentially private mean embedding of the real data. Given this differentially private embedding, the objective of the model’s generator does not have direct access to the data and can then be freely optimised. DP-NTK: Building on DP-MERF, Yang et al. [2024] proposes a model which utilises the features of an empirical neural tangent kernel rather than the random Fourier features used by DP-MERF. This modification enables the model to better capture complex, non-linear relationships in tabular data, improving the balance between privacy and utility over the original DP-MERF, without relying on any public tabular datasets.
@@ 4.6 Other Deep Learning Methods
Models adapted from other domains to tabular data synthesis. In (§3.1-3.5) and earlier in this section, we compared deep generative models originally proposed for tabular data synthesis. Here, we discuss seminal models first developed in other fields and later adapted for tabular data. WGAN: Originally designed for computer vision tasks, WGAN [Arjovsky et al., 2017] was introduced in a paper that analyses ways to measure the distance between the real and model distributions, arguing that weaker loss functions induce weaker model topologies, thus weak models for the real distribution. The authors compare the Wasserstein distance with traditional probability distance measures, including Kullback-Leibler (KL) divergence, Jensen-Shannon divergence, and total variation distance, showing that the Wasserstein distance has properties that would be better suited for learning distributions in low dimensional manifolds. They thus propose GAN models that use the Wasserstein distance as loss function. WGAN-GP: A popular extension of WGAN is WGAN-GP [Gulrajani et al., 2017], which added a new term in the loss that penalised the gradient norm of the discriminator w.r.t. the discriminator’s input, providing an alternative to the weight clipping that was leading to the generator retaining the real data points and adding Gaussian noise to fill the gaps. VEEGAN: VEEGAN [Srivastava et al., 2017] jointly trains a generator and a novel reconstructor network in order to avoid mode collapse. The reconstructor learns to map the real data distribution to Gaussian random noise, approximating the inverse of the generator and encouraging the generator to map the noise distribution back to the real data distribution.
4.7 Non-deep learning alternatives
While deep learning approaches for synthesising tabular data have gained significant traction, it is important to note that they are not always better than the statistical methods. Indeed, in certain scenarios, such methods can deliver competitive utility performance with lower computational costs and easier hyper-parameter tuning. Further, for use cases requiring strong privacy guarantees, non-deep learning models are often better at balancing the utility-privacy trade-off than their deep learning counterparts.
One example of a classical statistical model is Synthpop [Nowok et al., 2016], which is able to preserve multivariate relationships by modelling each tabular data feature conditional on the previously synthesised feature values. Generating synthetic data sequentially in this way avoids the training instability often seen with GAN-based models. More recently, Watson et al. [2023] introduced Adversarial Random Forests (ARF), which are tree-based generators able to combine the adversarial functionality of GANs with the robustness of tree ensembles to ultimately produce high-utility data. Yet another application domain where non-deep learning models excel is the privacy-preserving synthesis. For instance, rather than trying to approximate the full joint distribution of the real data, the Adaptive and Iterative Mechanism (AIM) [McKenna et al., 2022] directly optimises for pre-defined statistical queries and adds noise to a set of marginals of the data in order to satisfy differential privacy, and finally generates samples that best explain the marginals.
Finally, there are also works that bridge generative modelling and tree-based methods. For instance, Jolicoeur-Martineau et al. [2024] show that methods traditionally relying on deep learning, such as score-based diffusion and flow-matching, can be implemented using XGBoost instead of neural networks. This approach shows improved statistical fidelity and diversity compared to deep learning baselines such as TabDDPM Kotelnikov et al. [2023] particularly on small datasets. Building on this approach, Cresswell & Kim [2024] address the scalability limitations of these tree-based models, proposing efficient implementations that enable scaling to larger datasets with a larger number of features and samples.
@@ 5 Discussion and Future Directions
In this survey, we identified five key requirements for effective synthetic data deployment, alongside relevant evaluation procedures and deep generative models, while categorising the latter by model architecture.
@@
Related work. Previous surveys generally compare methods by the architecture-specific properties of the models. For example, both Borisov et al. [2024] and Davila R. et al. [2025] group methods mainly by their architectures, while also describing the necessary data transformation and regularisation techniques. [Lautrup et al., 2024] provide an overview of the models proposed in the field and the metrics used to evaluate them, while not taking this requirement-centric perspective which helps both newcomers and seasoned practitioners and researchers in the field to find the perfect model and evaluation protocol for their needs. Further, they exclude diffusion-based and LLM-based architectures. On the other hand, Wang et al. [2025] narrow the scope of their survey and focus specifically on methods for social and government sector applications, where balancing utility and privacy is the key concern.
Another related work is the survey by Jordon et al. [2022], which provides a valuable, high-level view of synthetic data across various modalities, including audio, image, video, time-series, and tabular data, without focusing specifically on the latter. However, with a primary emphasis on privacy, their survey treats requirements such as utility and fidelity as secondary goals. Further, rather than reviewing in detail methods for generating the data, the authors offer a high-level description of what synthetic data is and why it is useful to society, arguing, alongside Savage [2023], that a general public understanding of these methods is crucial given their potential to contribute to data democratisation. While Jordon et al. [2022] advocate for systematic frameworks, they categorise methods primarily by application (e.g., private data release, fairness). Further, regarding tabular data, their work relies on GAN-based approaches and does not cover recent advancements such as diffusion or transformer-based architectures. Finally, van Breugel et al. [2024] review methods for improving data fairness and reducing bias. However, their work focuses on biomedically-relevant data (including molecular, tabular, and imaging data) and surveys foundation models able to create such data via query-specific generation. Thus, rather than focusing on methods that generate full datasets intended to replace or augment real data (i.e., the focus of this survey), they review methods which generate single samples conditioned on specific inputs (e.g., an answer in the form of an image to a medical question).
In contrast to these works, our survey distinguishes itself by providing a comprehensive overview of the requirements specific to synthesising tabular data and their complex interdependencies. To this end, we offer a practical, high-level roadmap (Figure 1) to guide users in selecting appropriate models. We support this roadmap with granular decision-making tools, including: (i) a structured technical taxonomy offering architectural details (Section 4), (ii) method-related specifics such as mixed data type support, inference time, and scalability to high-dimensional data (as part of Table 1), and (iii) an overview on which requirements are addressed by each method (Figure 3). The latter is particularly valuable for refining model selection based on requirements that are too intertwined to be effectively captured in the initial high-level roadmap. Furthermore, we support this user-centric perspective by detailing evaluation protocols and metrics for each requirement (Section 3). Overall, this structure positions our survey as a resource for both newcomers and specialists to select and evaluate models tailored to their specific application scenarios.
Future Directions. Tabular data generation is a rapidly evolving field still in its early stages. As illustrated in Figure 3, many early approaches have prioritised individual objectives—typically utility or privacy—leading to highly specialised models that lack broad applicability. Note that the categorisation in Figure 3 should be interpreted as a reflection of each method’s focus, rather than an exclusive capability. As discussed in Section 3.6, these requirements are deeply interconnected: they often overlap or trade-off against one another. For example, the methods associated solely with the privacy requirement are positioned as such because their core contribution is a privacy-preserving mechanism. However, this does not implicitly suggest they lack utility (indeed, a synthesiser with very low utility would be useless in any downstream task), but rather that their innovations are driven by primarily addressing the privacy requirement. Similarly, the methods associated to the fidelity requirement focus on approximating the real data distribution, which often sets the foundation for achieving a high utility.
This fragmentation highlights a key challenge: the need for models that can balance multiple, often competing, requirements. Notably, alignment has historically received limited attention but is now emerging as a crucial factor for generating realistic data. With the increasing deployment of generative models in real-world applications and the growing prominence of neurosymbolic AI and safe AI [d’Avila Garcez et al., 2019; Giunchiglia et al., 2022], we anticipate that alignment will soon become a fundamental requirement in the field. Emerging approaches are already tackling this challenge. For instance, training-based methods like Stoian & Giunchiglia [2025] introduce new layers to guarantee constraint satisfaction and improve utility, while alternative directions focus on developing methods that do not require retraining, such as the work in Scassola et al. [2025b], which proposes a zero-shot framework utilising soft constraints to approximate the conditional distribution. Moreover, the Figure reveals a striking gap: no existing model successfully integrates all the core requirements we consider. As the field matures, we expect new requirements will emerge, and the hybridisation of techniques will drive the development of more versatile models. Bridging these gaps will be essential for advancing the field and enabling broader, more reliable applications of tabular data generation, potentially expanding to complex multi-table relational data settings as recently explored by Scassola et al. [2025a].
Broader Impact Statement
This survey captures current trends in deep generative modelling for tabular data synthesis, offers a new perspective on what makes synthetic data practically useful, and outlines key steps toward achieving such a goal. We emphasise that, while existing methods address different requirements (e.g., alignment, utility, privacy, and fidelity), further progress toward developing approaches that simultaneously meet multiple requirements will increase their impact in real-world applications. In particular, we highlight the importance of incorporating domain knowledge to enhance the quality of synthetic data and encourage its integration in future deep generative modelling approaches for synthesising tabular data. It is important, however, to ensure that such background knowledge does not embed user-specific biases or characteristics, as this could raise privacy concerns. Provided this condition is met, injecting domain knowledge into generative models does not pose any direct negative impact.
Acknowledgements